OpenStudy (anonymous):

a radioactive isotope has a half-life of approximately 1850 years. How many years to the nearest year would be required for a given amount of this isotope to decay 40%?

OpenStudy (anonymous):

= 40% * 1850 / 50% = 0.4 * 1850 / 0.5 = 1480 years